プレゼンテーション技法 (D) 2 単位 (選択)

Presentation Method (D)  先端技術科学教育部

【授業目的】発表 (学術雑誌，口頭発表，ポスター発表，新聞等メディア) することで，研究はその成果が生きる．様々な発表手段のうち，学会での発表は討論，情報交換，アイデアの展開とそれに続く研究の飛躍を図るため重要である．本講義は，学会において効率よく最大限のアピールができるプレゼンテーションの方法を学ぶ．

【授業概要】国際学会で学生が自ら発表するように指導を受け，それを実行した

場合に単位を認定しようとする授業科目である．【授業形式】 [形態] 【キーワード】 [キーワード] 【先行科目】 [先行科目] 【関連科目】 [関連科目] 【履修要件】 [要件] 【履修上の注意】 [注意] 【到達目標】学会，会議における発表の知識 · 経験を有する．【授業計画】 [計画] 【成績評価基準】学会発表の準備過程におけるポートフォリオ，学会発表報告書【教科書】 [教科書] 【参考書】 [参考資料] 【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216873 【対象学生】開講コース学生のみ履修可能【連絡先】 [連絡先] 【備考】 90 時間分の学修記録 (ポートフォリオ) を提出する．詳細については，指

導教員に相談，指示を仰ぐこと．

Presentation Method (D) 2 units (selection)

Graduate School of Advanced Technology and Science

Target⟩ Poster and oral presentations, and journal publications are common means of appealing the research work in various scientific and technological disciplines. It’s undoubtedly important for activating the research work. The lecture is intending for students to learn most efficient way to appeal their findings and discussions in scientific community.

Outline⟩ Credits are acknowledged to the students who are supervised to give

conference speech by themselves and do actually the lectures. Style⟩ [形態] �������⟩ [キーワード] Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Requirement⟩ [要件] Notice⟩ [注意] Goal⟩ The lecture is aiming at having knowledge and experience of the presentaion

in international conferences. Schedule⟩ [計画] Evaluation Criteria⟩ Students’ achievement is evaluated by portfolios submitted

showing the progress in preparation of international conferences as well as by presentation records.

Textbook⟩ [教科書] Reference⟩ [参考資料] Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216873

Student⟩ Able to be taken by only specified class(es) Contact⟩ [連絡先] Note⟩ A possible student should submit portfolios which are evidenced for his/her

90 hours’ extensive work. Students are advised to ask details about this lecture to their supervisers.